What is the wavelength of electromagnetic energy that could excite an electron in a hydrogen atom from n=1 to n=3 energy levels? ΔE=-2.18 x 10-18 J ·(1/(n_f^2 )-1/(n_i^2 ))
Answer #1

Rydberg equation: E=-2.18 x 10-18 J E=-2.18-10-{} } E =-2.18*10*287 =-1.94 x 10-18 J Therefore, E=he 6.625x10-34 J-s x 3.00